OpenStudy (hlares):

Question: "For which n does the expression 1+2+...+(n-1) congruent to 0(mod n) hold?" If I am reading this correctly, for the expression 0 is the least positive residue of (1+2+...+(n-1))(mod n). However, if that is correct, I am not sure where to go from there. Would n be 0 or 1 then? Or something else entirely?

OpenStudy (freckles):

let's look at some examples let's say n is 1... then you have 0 congruent to 0 mod 1 let's say n is 2 then you have 0+1 congruent to 0 mod 2 which is false let's say n is 3 what happens?

OpenStudy (hlares):

Then if n=3, it would be 1+2=3 congruent to 0mod3, which would be true then, aye?

OpenStudy (freckles):

yes so let's look at the more general thing you have 0+1+2+3+...+(n-1) or 1+2+3+...+(n-1) or 1+(n-1) + 2+(n-2) + 3+(n-3) +..... now I can pair the terms of the sum up if there are an even amount of terms by the way you do know that k+(n-k) is n right? so each of those rows I wrote have sum n and what is n mod n

OpenStudy (freckles):

for example n=4 gives 1+2+3 which doesn't have a odd amount of terms you have (1+3)+2 for n=5 you have 1+2+3+4 you have an even amount of terms so we have (1+4)+(2+3) for n=6 you have 1+2+3+4+5 or (1+5)+(2+4)+3 ...

OpenStudy (freckles):

do you know how to evaluate k+(n-k) mod n?

OpenStudy (hlares):

I do not off the top of my head. I am looking back through my text right now to see if it was mentioned and if so where. As for n(mod n), that would be 0 though, aye?

OpenStudy (freckles):

right and k+(n-k) is n so k+(n-k) mod n is 0

OpenStudy (hlares):

Ah okay! And if I am understanding pairing up the terms correctly: that is being done to see if the sum is divisible by n, right? For example, for n=6, that would be false because it has an extra 3, aye?

OpenStudy (freckles):

right and 3 mod 6 isn't 0

OpenStudy (freckles):

6+6+3=3 but 3 mod 6 isn't 0

OpenStudy (freckles):

if you have an even amount of terms in your sum you can pair them up and such away you have n+n+n+n+...+n mod n

OpenStudy (freckles):

I think I also have another way to look at it too after you get this way

OpenStudy (hlares):

Okay, that makes sense! We did not cover pairing them up like that at all, so that's really useful to know, thank you. And if I am noticing the pattern correctly here, the only n's for which the statement hold thus far are for odd integers as the even ones result in false statements, right?

OpenStudy (freckles):

right

OpenStudy (freckles):

another way: now \[1+2+3+ \cdots +(n-1)=\frac{n(n-1)}{2} \text{ do you remember this formula }\]

OpenStudy (freckles):

think back to calculus 1 days

OpenStudy (freckles):

if that isn't too far back

OpenStudy (freckles):

\[\sum_{i=1}^{n}i=\frac{n(n+1)}{2}\]

OpenStudy (freckles):

except we have to n-1

OpenStudy (hlares):

Calculus 1 was awhile ago, but that was covered under a section on mathematical induction in this class recently.

OpenStudy (freckles):

\[\sum_{i=1}^{n-1}i=\frac{(n-1)n}{2} \text{ or } \frac{n(n-1)}{2}\]

OpenStudy (freckles):

oh cool

OpenStudy (freckles):

so we can solve this equation \[\frac{n(n-1)}{2} \equiv 0 \mod n \]

OpenStudy (freckles):

now if n-1 is even then we can cancel out the two's and we are left with \[\frac{n(2a)}{2} =n a \equiv 0 a \equiv 0 (\mod n)\]

OpenStudy (freckles):

if n-1 is even then n is odd

OpenStudy (freckles):

if n-1 is odd we can't cancel out the 2 on the bottom but n would be even and n would get destroyed by that division by 2 if you know what I mean so we wouldn't get 0 when modding it by n

OpenStudy (freckles):

you can write the fraction like this: \[n \frac{n-1}{2} \equiv 0 \mod n \text{ As long as } \frac{n-1}{2} \in \mathbb{Z} \]

OpenStudy (hlares):

Aye, I do. That makes a lot of sense. Thank you for explaining it so well; your explanation was a lot more understandable than the section of the text relating to this problem.

OpenStudy (freckles):

np

OpenStudy (freckles):

you know when (n-1)/2 is an integer right? I kind of already said but just want to be sure you know

OpenStudy (hlares):

Aye, I do! Can only deal with integers with congruences, right?

OpenStudy (freckles):

well I think you can actually use other numbers but in number theory yeah I think you only use integers that is all I ever seen in number theory n is a positive integer and... for (n-1)/2 to be an integer n-1 would have to be even because that is the only sorta of number to be divisible by 2 and if n-1 is even then n is odd since n and n-1 are consecutive integers

OpenStudy (hlares):

Ah okay. This is all relating to brief stint into elementary number theory for this class, so all we have dealt with with problems is either all integers or positive integers/natural numbers, which is why I presumed that. And okay, that makes sense!

OpenStudy (freckles):

anyways glad I can help

OpenStudy (hlares):

Thank you very much for the assistance, I greatly appreciate it!